hi. Could you help me understand standard deviation? Here is my problem:
Consider the following data set: 28, 34, 41, 19, 17, 23.
a) Compute the mean and standard deviation of the data set
b) Suppose each value is mulitplied by two and then 5 is added to each of the values after each is multiplied by two. Compute the mean and standard deviation of this new data set.
Thanks for your help!!!

Answer by venugopalramana(3286)

Could you help me understand standard deviation? Here is my problem:
Consider the following data set: 28, 34, 41, 19, 17, 23.
a) Compute the mean and standard deviation of the data set
b) Suppose each value is mulitplied by two and then 5 is added to each of the values after each is multiplied by two. Compute the mean and standard deviation of this new data set.
Thanks for your help!!!
......... DEVIATION SQUARE OF
S.NO..... X WRT.MEAN DEVIATION
1.......... 28 1 .............1
2....... 34 7 ............49
3....... 41 14 ...........196
4....... 19 -8 ............64
5....... 17 -10 ............100
6....... 23 -4 .............16
---------------------------------------------------
TOTAL... 162 0 .............426
MEAN=162/6=... 27...... MEAN.OF SQUARES =426/6=71
........OF DEVIATION
R.M.S.DEVN.=SQRT(71)= 8.43
OR STD.DEVN.

DEVIATION SQUARE OF
S.NO. X WRT.MEAN DEVIATION
1 61 34...... 1156
2 73 46......... 2116
3 87 60...... 3600
4 43 16....... 256
5 39 12...... 144
6 51 24...... 576
TOTAL 354 192..... 7848
MEAN=354/6= 59...... MEAN.OF SQUARES
................OF DEVNS.
.........=7848/6=1308

R.M.S.DEVN. = SQRT(1308)=36.16628264
OR STD.DEVN.